# Aspen Open Jets: Unlocking LHC Data for Foundation Models in Particle Physics

### v1

@@ III.2 Performance metrics
    The probability densities $P(x)$ and $Q(x)$ for the observable $\mathcal{O}$ are obtained from the normalized histogram bin counts and ${\rm KL}^{\mathcal{O}}(P||Q)$ is computed using scipy.stats.entropy Virtanen et al. (2020). Note that ${\rm KL}^{\mathcal{O}}(P||Q)$ is antisymmetric. In our results, we take $P$ to be the generated distribution and $Q$ to be the reference JetClass distribution.
@@
    $W^{\mathcal{O}}_{1}(P,Q)=\min_{\pi\in\Pi}\,\sum_{x,y}\absolutevalue{x-y}\ \pi(x,y)\,.$ (2)
    Here $\Pi(P,Q)$ is the space of all joint ‘coupling’ distributions whose marginals are the observable $\mathcal{O}$ distributions $P$ and $Q$. This quantity is computed with scipy.stats.wasserstein_distance Virtanen et al. (2020). Unlike the ${\rm KL}^{\mathcal{O}}(P||Q)$ divergence, $W^{\mathcal{O}}_{1}(P,Q)$ is a symmetric similarity metric based on optimal transport, i.e. the optimal cost plan that morphs $P$ into $Q$.
@@ III.3 Results
Overall, we find that the pre-trained models are able to achieve a performance gain when using a fraction of the total fine-tuning data. This can be seen in the third and last columns of fig. 3 and fig. 4, where the fine-tuned models (blue lines) typically achieve better [^6] However, there are instances where models trained from scratch perform equally well on a single metric within uncertainties. For example, this occurs with ${\rm KL}^{p_{T}}$ for training sample size $D=10^{3}$. This behavior can be attributed to the definition of KL divergence, which penalizes high-probability regions in the generated data where the true probability distribution is low. Therefore, it is crucial to evaluate performance using multiple metrics to gain a comprehensive understanding. metric scores than the models trained from scratch (red lines). This is apparent for the KL divergence and Wasserstein-1 metrics across different training sample sizes below $1\,\rm M$ jets. However, as the available training data exceeds $1\,\rm M$ jets, we find that the fine-tuning begins to saturate, and the fine-tuned and from-scratch models perform similarly within uncertainties. In some features, the fine-tuned models even exhibit a slight decrease in performance compared to those trained entirely from scratch, indicating a diminishing return from pre-training at larger dataset scales. The reasons for this are unclear and would require further study. One possibility is that the AOJ tokenization is suboptimal for describing JetClass and this is exposed with enough JetClass training data. Another possibility is that the pre-training itself can harm model performance by locking in the weights to a suboptimal part of parameter space, a phenomenon that has been observed previously in the ML literature Hernandez et al. (2021) and was termed “ossification”.
@@ III.4 Power-law scaling
Finally, we observe that the performance metrics for the fine-tuned models, when not in the saturated regime, follow a power-law scaling of the form $D^{-\gamma}$ with respect to the training sample size $D$. Neural scaling laws, such as these, have attracted broad interest and have been analyzed extensively in various settings in recent years, see Villalobos (2023) for a nice overview of the ML literature and original references. In particular, they have been observed in the context of jet classification Batson and Kahn (2023). We find that the exponents for a wide range of high-level observables are consistently within the range $0.450.2<\gamma<$ 0.45 (see dashed lines in fig. 3 and 4). This scaling behavior holds for both the KL and Wasserstein-1 metrics and applies to both QCD and top-quark, suggesting a form of universality. The only exception to this scaling is the particle multiplicity distribution, which shows a more irregular dependence on the training sample size. Overall, this demonstrates that the generative performance scales predictably and uniformly with the training sample size across most high-level observables, enabling a smooth and consistent transfer learning between the AOJ and JetClass datasets.
In contrast, the randomly initialized models trained from scratch do not satisfy such universal scaling across most physical observables. The reason behind this may be that these operate in the small-data regime Hestness et al. (2017), where their performance is limited by the lack of sufficient training data. In this regime, models can only perform marginally better than random guessing, as each additional sample provides limited new information for learning. The breaking of scaling is particularly evident for the jet substructure features, such as $\tau_{21}$ for QCD and $\tau_{32}$ for tops, and to a lesser extent for the jet mass, where the performance metrics exhibit irregular and dataset-dependent scaling as the training sample size increases. Only after the model has seen sufficient training data, approximately around $1\,\rm M$ jets, does the training stabilize, resulting in reliable jet substructure and jet mass modeling. This discrepancy highlights the advantage of pre-training in achieving a more robust and systematic improvement in generative performance.
@@ IV Conclusion
We are releasing the AspenOpenJets (AOJ) dataset, which comprises 180 million jets from CMS open data, in a standard machine learning-friendly format. This dataset provides an excellent resource for exploring foundation models and other machine learning techniques that use large, unlabelled datasets for pre-training.
In addition, we present the first jet-based foundation model in high-energy physics trained on real collider data from the AOJ dataset. By using this rich dataset to pre-train a foundation model, we show that fine-tuning for specific tasks - such as generating jets of different types in different kinematic regions - yields significant performance improvements over models trained from scratch. In particular, the generative performance of our foundation model scales predictably with training sample size, facilitating effective transfer learning. This demonstrates the power of foundation models to provide improved performance even over significant domain shifts.
@@ Appendix A CMS Samples
MC samples provided by CMS were used to estimate the composition of the AOJ sample. CMS uses a variety of generators for the simulation different physics processes, including MADGRAPH5_aMC@NLO Alwall et al. (2014), PYTHIA Sjöstrand et al. (2015), and POWHEG Nason (2004); Frixione et al. (2007); Alioli et al. (2010). Each sample uses PYTHIA with the underlying event tune CP5 Sirunyan et al. (2020c) to simulate the parton shower and hadronization. All samples use the NNLO NNPDF 3.1 parton distribution functions Ball et al. (2010, 2015, 2017). Some physics processes are split into several samples covering different kinematic regions. Each sample is normalized to the appropriate cross section.

### v2

@@ III.2 Performance metrics
    The probability densities $P(x)$ and $Q(x)$ for the observable $\mathcal{O}$ are obtained from the normalized histogram bin counts and ${\rm KL}^{\mathcal{O}}(P||Q)$ is computed using scipy.stats.entropy 2020SciPy-NMeth. Note that ${\rm KL}^{\mathcal{O}}(P||Q)$ is antisymmetric. In our results, we take $P$ to be the generated distribution and $Q$ to be the reference JetClass distribution.
@@
[…]
    Here $\Pi(P,Q)$ is the space of all joint ‘coupling’ distributions whose marginals are the observable $\mathcal{O}$ distributions $P$ and $Q$. This quantity is computed with scipy.stats.wasserstein_distance 2020SciPy-NMeth. Unlike the ${\rm KL}^{\mathcal{O}}(P||Q)$ divergence, $W^{\mathcal{O}}_{1}(P,Q)$ is a symmetric similarity metric based on optimal transport, i.e. the optimal cost plan that morphs $P$ into $Q$.
While further detailed performance evaluations such as classifier tests Krause:2021ilc; Kansal:2022spb; Das:2023ktd could also be envisioned, this paper is first and foremost demonstrating a use case for the AOJ dataset in the context of foundation models. As such, we are rather concerned with the overall qualitative behavior, for which the 1D distributions suffice.
@@ III.3 Results
Overall, we find that the pre-trained models are able to achieve a performance gain when using a fraction of the total fine-tuning data. This can be seen in the third and last columns of fig. 3 and fig. 4, where the fine-tuned models (blue lines) typically achieve better [^7] However, there are instances where models trained from scratch perform equally well on a single metric within uncertainties. For example, this occurs with ${\rm KL}^{p_{T}}$ for training sample size $D=10^{3}$. This behavior can be attributed to the definition of KL divergence, which penalizes high-probability regions in the generated data where the true probability distribution is low. Therefore, it is crucial to evaluate performance using multiple metrics to gain a comprehensive understanding. metric scores than the models trained from scratch (red lines). This is apparent for the KL divergence and Wasserstein-1 metrics across different training sample sizes below $1\,\rm M$ jets. However, as the available training data exceeds $1\,\rm M$ jets, we find that the fine-tuning begins to saturate, and the fine-tuned and from-scratch models perform similarly within uncertainties. In some features, the fine-tuned models even exhibit a slight decrease in performance compared to those trained entirely from scratch, indicating a diminishing return from pre-training at larger dataset scales. The reasons for this are unclear and would require further study. One possibility is that the AOJ tokenization is suboptimal for describing JetClass and this is exposed with enough JetClass training data. Another possibility is that the pre-training itself can harm model performance by locking in the weights to a suboptimal part of parameter space, a phenomenon that has been observed previously in the ML literature DBLP:journals/corr/abs-2102-01293 and was termed “ossification”.
@@ III.4 Power-law scaling
Finally, we observe that the performance metrics for the fine-tuned models, when not in the saturated regime, seem to follow a power-law scaling of the form $D^{-\gamma}$ with respect to the training sample size $D$. Neural scaling laws, such as these, have attracted broad interest and have been analyzed extensively in various settings in recent years, see epoch2023scalinglawsliteraturereview for a nice overview of the ML literature and original references. In particular, they have been observed in the context of jet classification Batson:2023ohn. We find that the exponents for a wide range of high-level observables are consistently within the range $0.2<\gamma<0.45$ (see dashed lines in fig. 3 and 4). This scaling behavior holds for both the KL and Wasserstein-1 metrics and applies to both QCD and top-quark, suggesting a form of universality. The only exception to this scaling is the particle multiplicity distribution, which shows a more irregular dependence on the training sample size. Previous scaling law behavior has been observed as a function of the loss of the model, not on such hand-crafted metrics. However, due to the different tokenizers used in the from-scratch and the AOJ fine-tuned model, the losses for the generation objective are not directly comparable, making statements of true scaling comparisons difficult. We leave further investigation of this apparent scaling behavior to future work.
In contrast, the randomly initialized models trained from scratch are not seen to follow such universal scaling across most physical observables. The reason behind this may be that these operate in the small-data regime hestness2017deep, where their performance is limited by the lack of sufficient training data. In this regime, models can only perform marginally better than random guessing, as each additional sample provides limited new information for learning. The breaking of scaling is particularly evident for the jet substructure features, such as $\tau_{21}$ for QCD and $\tau_{32}$ for tops, and to a lesser extent for the jet mass, where the performance metrics exhibit irregular and dataset-dependent scaling as the training sample size increases. Only after the model has seen sufficient training data, approximately around $1\,\rm M$ jets, does the training stabilize, resulting in reliable jet substructure and jet mass modeling. While further studies into these scaling laws need to be performed, the observed discrepancy highlights the possible advantage of pre-training in achieving a more robust and systematic improvement in generative performance.
@@ IV Conclusion
We are releasing the AspenOpenJets (AOJ) dataset, which comprises 178 million jets from CMS open data, in a standard machine learning-friendly format. This dataset provides an excellent resource for exploring foundation models and other machine learning techniques that use large, unlabelled datasets for pre-training.
In addition, we present the first jet-based foundation model in high-energy physics trained on real collider data from the AOJ dataset. By using this rich dataset to pre-train a foundation model, we show that fine-tuning for specific tasks - such as generating jets of different types in different kinematic regions - yields significant performance improvements over models trained from scratch, and boosts the achievable downstream performance on small labeled datasets. In particular, the generative performance of our foundation model scales predictably with training sample size, facilitating effective transfer learning. This demonstrates the power of foundation models to provide improved performance even over significant domain shifts.
@@ Appendix A CMS Samples
MC samples provided by CMS were used to estimate the composition of the AOJ sample. CMS uses a variety of generators for the simulation different physics processes, including MADGRAPH5_aMC@NLO Alwall:2014hca, PYTHIA Sjostrand:2014zea, and POWHEG Nason:2004rx; Frixione:2007nw; Alioli:2010xd. Each sample uses PYTHIA with the underlying event tune CP5 CMS:2019csb to simulate the parton shower and hadronization. All samples use the NNLO NNPDF 3.1 parton distribution functions Ball:2010de; NNPDF:2014otw; NNPDF:2017mvq. Some physics processes are split into several samples covering different kinematic regions. Each sample is normalized to the appropriate cross section.